OpenStudy (anonymous):

The Maryland Monkeys won six less than twice as many soccer games as The Pennsylvania Pigs. If the teams won a total of 48 games,how many games did each win ?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

In alegebraric solution !!

OpenStudy (anonymous):

ok. whats the answer choices?

OpenStudy (anonymous):

The answer is 18 games and 30 but i just need to do the problem but i dont know what to do

OpenStudy (anonymous):

You divide 48 into two .. Then you will have 24 ... Is 24 any of you answer choice?

OpenStudy (anonymous):

No just 18 and 30

OpenStudy (anonymous):

Is in algebraric solution !!

OpenStudy (anonymous):

dang i though i knew my bad

OpenStudy (anonymous):

Oh ok but anyways thanks :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Can somebody help me with this problem ????

OpenStudy (anonymous):

Trying to figure the problem for you and THANK YOU FOR FANNING ME @Emily55

OpenStudy (anonymous):

You got the answer in the back of the book right and have no idea how to do it

OpenStudy (anonymous):

Yea i dont have any idea

OpenStudy (anonymous):

Hello @Chibiterasu Its in your book @Emily55 the problem

OpenStudy (♪chibiterasu):

Well we first set up a basic equation where m will represent the Maryland Monkeys and p will represent the Pennsylvania Pigs m = 2p - 6 m + p = 48

OpenStudy (♪chibiterasu):

According to a few recourses, the Monkeys got 30 and the Pigs got 18.

OpenStudy (♪chibiterasu):

Not so sure about solving though, it's been a while since I've done problems like these.

OpenStudy (anonymous):

@Emily55

OpenStudy (anonymous):

you still there?

OpenStudy (mathlegend):

So I think we should do this... they won six less than the other team... x-6

OpenStudy (mathlegend):

Then since it says twice as many... 2(x-6) = 48 (since this is there total)

OpenStudy (mathlegend):

@Emily55 can you solve this for 'x'?

OpenStudy (mathlegend):

Well does anyone else want to know how you get the answer?

OpenStudy (anonymous):

I'm solve the equation you gave @Emily55

OpenStudy (anonymous):

30 is what I got @MathLegend but see I know how to solve it than struggle on the part where you set up the equation

OpenStudy (mathlegend):

Yeah, I always have to read word problems at least 2-3 times myself. :) Just takes patience. Then note 30 is for the team of Maryland... and since the total won was 48... you subtract... 48-30 = 18 (which is the other team... I forget their name)

OpenStudy (anonymous):

I did it like 2x-6=48 and 6x-2=48 and realize I didn't get 30 or 18 Maryland Monkeys and Pennsylvania Pigs.

OpenStudy (anonymous):

YOU DESERVE A MEDAL @MathLegend

OpenStudy (mathlegend):

Thank You!

OpenStudy (anonymous):

You REALLY DESERVE IT because you solve it. I thought I could get it but no I couldn't

OpenStudy (anonymous):

THANKS GYS FORM HELPING ME